The lives of 100 batteries of Brand D and 100 batteries of Brand E were determined. The results ane Exercises 11-15 summarized in the histograms below. Brand D 40- 30- 20- 10- 105 165 85 125 95 115 135 145 155 ife (hours) nol Brand E 35 30 STE 25 20 15 10 95 65 115 Life (hours) 155 105 135 145 125 isuanbeu kousnbeu
11 Estimate the mean battery life for Brand D. (Do not do any calculations.) are 12 Estimate the mean battery life for Brand E. (Do not do any calculations.) Which of Brands D and E shows the greater variability in battery lives? Do you think the two brands are roughly the same in this regard? 14. Estimate the largest deviation from the mean for Brand D. 15. What would you consider a tvpical deviation from the mean for Brand D? Pace /28 Page | 29
Answer #1

(11) answer is 110

since the frequency of the class 105-115, is highest and looking almost symmetrical about this class, so mean would be mid value of 105 and 115 i.e. (105+115)/2=110

(12)answer is 130

since the frequency of the class 125-135, is highest and looking almost symmetrical about this class, so mean would be mid value of 125 and 135 i.e. (125+135)/2=130

(13) Brand D showed more variability, as the height of the class in the brand D is more heterogeneous than brand E

(14) largest deviation=largest observation-mean=140-110=30

(15) it would be zero, as it is almost symmetrical about the mean value 110 (i.e. class 105-115)